In one day, a store sold 22 lamps. The larger lamps sell for $20 and the smaller lamps sell for $10. The total sales for the day were $320. How many of each lamp was sold ?
Answer by mananth(16949)

larger lamps x numbers -------------$20
smaller lamps --- 22-x -------------$10
..
sales = $320
20x+10(22-x)=320
20x+220-10x=320
10x+220=320
-220
10x= 100
/10
x= 10 -------larger lamps
balance 12 smaller lamps
...
m.ananth@hotmail.ca